11. Show that if A and B are sets, then
(a) A − B = A ∩ B
(b) (A ∩ B) ∪ (A ∩ B) = A
1
Expert's answer
2020-07-27T18:06:35-0400

a) "A-B=A\cap B^C"

We must show "A-B\sube A\cap B^C" and "A\cap B^C\sube A-B"

Show that "A-B\sube A\cap B^C"

Let "x\in A-B." By definition of set difference, "x\in A" and "x\notin B." By definition of complement, "x\notin B" implies that "x\in B^C." Hence, it is true that both, "x\in A" and "x\in B^C." By definition of intersection, "x\in A\cap B^C."

Show that "A\cap B^C\sube A-B."

Let "x\in A\cap B^C." By definition of intersection, "x\in A" and "x\in B^C." By definition of complement, "x\in B^C" implies that "x\notin B." Hence, "x\in A" and "x\notin B." By definition of set difference, "x\in A-B."

Thus, "A-B=A\cap B^C."


b) "(A-B)\cup(A\cap B)=A"

Let "x\in A." There are two cases, "x\in A" and "x\notin B" or "x\in A" and "x\in B."

In the first case "x\in A, x\notin B," so by definition set difference, "x\in A-B."

In the second case "x\in A" and "x\in B," so by definition of intersection "x\in A\cap B."

By definition of union "x\in (A-B)\cup(A\cap B)."

Thus if "x\in A," then "x\in (A-B)\cup(A\cap B)."


Let "x\in (A-B)\cup(A\cap B)." This means that either "x\in A-B" or "x\in A\cap B."

In the first case "x\in A, x\notin B," in the second case "x\in A" and "x\in B."

Then in either case "x\in A."

Two sets are equal, since they have the same elements.Therefore


"(A-B)\cup(A\cap B)=A"